OpenStudy (anonymous):

subtract x^2-1/x^2-x-2 - x-1/x-2

OpenStudy (anonymous):

factorize \[x^2-1\] and \[x^2-x-2\]

OpenStudy (anonymous):

so it would be x-1/x-2??

OpenStudy (anonymous):

the answer i mean

OpenStudy (anonymous):

no \[x^2-1= (x-1)(x+1)\] and \[x^2-x-2 = x^2 - 2x + x -2 = x(x-2) + 1(x-2)= (x+1)(x-2)\] so your problem is \[{{x^2-1}\over{x^2-x-2}}-{{x-1}\over{x-2}}= {{(x-1)(x+1)}\over{(x+1)(x-2)}} - {{x-1}\over{x-2}} = {{x-1}\over{x-2}}-{{x-1}\over{x-2}}=0\]

OpenStudy (anonymous):

so the answer is 0

OpenStudy (anonymous):

oh ok it got it!

OpenStudy (anonymous):

yeah buddy you were right

OpenStudy (anonymous):

according to the SFPD we are Locating the location of hayesxelite for improper use and bad language on the internet the police will be at your house in less than 24 hours for questioning, thank you SFPD San Francisco Police Department.

OpenStudy (anonymous):

CAN YOU send those guys to my house toooo???? i want a piggy bak ride ??? ;) @stupid12345678910